Ice has a density of 0.92 g/cm3 and seawater has a density of 1.04 g/cm3

. Determine the total

volume of a floating iceberg if 30 m3

is above the seawater


1
Expert's answer
2021-05-09T15:13:58-0400

We have given that,

Density of ice = "0.92g\/cm^3"

Density of Sea water = "1.04 g\/cm^3"

Volume of iceberg above sea water = "30m^3"

Let v be the volume of iceberg immersed in water


As per the law of floatation, "\dfrac{v}{V} = \dfrac{Density \hspace{3mm}of\hspace{3mm} ice }{Density\hspace{3mm} of \hspace{2mm} sea \hspace{3mm}water}"


"\dfrac{v}{V} = \dfrac{0.92}{1.04} = 0.88"


"v = 0.88V"


Volume of the iceberg over the surface of sea water is given by:


"V(1-0.88) = 30"


"v = \dfrac{300}{0.12}"


"v = 250m^3"